OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 64. What is the length of the altitude? A) 4 B) 8 C) 16 D) 32

jhonyy9 (jhonyy9):

so from the postulate of height we know that the lenght of height squared is equal to the sum of their proiects squared

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

so hence the lenght of altitude will be squarroot(16^2 +64^2)

jhonyy9 (jhonyy9):

do you know the height item inside a triangle ?

OpenStudy (anonymous):

no.

jhonyy9 (jhonyy9):

check it on wikiipedia and will be secure that you will can understanding it sure

jhonyy9 (jhonyy9):

so i think that there will be wrote for you understandably

jhonyy9 (jhonyy9):

sorry for my english

jhonyy9 (jhonyy9):

so you have checked it there please ?

OpenStudy (anonymous):

i don't understand it .

OpenStudy (anonymous):

my mom just explained it to me . i'm good now :) thanks!!

jhonyy9 (jhonyy9):

ok good luck bye

Directrix (directrix):

Theorem: If the altitude is drawn to the hypotenuse of a right triangle, the length of the altitude is the geometric mean between the lengths of the segments of the hypotenuse. If a represents the length of the altitude, then the theorem says that 16 is to a as a is to 64. 16/a = a/64 a^2 = 16*64 a = 4* 8 a = 32 -----> Answer -->D) 32

OpenStudy (anonymous):

Thanks , I would've gotten that wrong :)